QUESTION 2 Youth Corporation has an activity-based costing system with three activity cost pools-Processing Setting Up, and Other. The companys overhead costs consist of equipment depreciation and indirect labor and are allocated to the cost pools in proportion to the activity cost pools consumption of resources. Equipment depreciation totals $72,000 and indirect labor totals $8,000. Data concerning the distribution of resource consumption across activity cost pools appear below Equipment depreciation020 Indirect labor Processing Setting Upother 0.40 0.50 0.40 Required: Assign overhead costs to activity cost pools using activity-based costing.
Answer #1
i ii iii iv v=iv*i vi=iv*ii vii=iv*iii
cost pool Allocation of overhead cost
Processing Setting up Other Total cost Processing Setting up Other
Equipment depreciation 0.2 0.4 0.4 72000 14400 28800 28800
Indirect labor 0.2 0.3 0.5 8000 1600 2400 4000